### v1

# A Comprehensive Multi-Vocal Empirical Study of ML Cloud Service Misuses

Abstract.
Machine Learning (ML) models are widely used across various domains, including medical diagnostics and autonomous driving. To support this growth, cloud providers offer ML services to ease the integration of ML components in software systems. The evolving business requirements and the popularity of ML services have led practitioners of all skill levels to implement, and maintain ML service-based systems. However, they may not always adhere to optimal design and usage practices for ML cloud services, resulting in common misuse which could significantly degrade the quality of ML service-based systems and adversely affect their maintenance and evolution. Though much research has been conducted on ML service misuse, a consistent terminology and specification for these misuses remain absent. We therefore conduct in this paper a comprehensive, multi-vocal empirical study exploring the prevalence of ML cloud service misuses in practice. We propose a catalog of 20 ML cloud service misuses, most of which have not been studied in prior research. To achieve this, we conducted a) a systematic literature review of studies on ML misuses, b) a gray literature review of the official documentation provided by major cloud providers, c) an empirical analysis of a curated set of 377 ML service-based systems on GitHub, and d) a survey with 50 ML practitioners. Our results show that ML service misuses are common in both open-source projects and industry, often stemming from a lack of understanding of service capabilities, and insufficient documentation. This emphasizes the importance of ongoing education in best practices for ML services, which is the focus of this paper, while also highlighting the need for tools to automatically detect and refactor ML misuses.
$c$s: ML Service, Misuse, Software Quality
1. Introduction
Major breakthroughs in Artificial Intelligence (AI) in general, and Machine Learning (ML) in particular, have brought huge success to many application domains such as image recognition, security, and autonomous driving (grigorescu2020survey, ; dong2021survey, ).
To support this technological hype, several major cloud providers like Amazon, Google, and Microsoft, have been pivotal in providing ML practitioners with various ML cloud infrastructure capabilities and specialized services necessary for performing ML-related tasks (wan2021machine, ). The availability of ML cloud services, which do not require designing or building ML models from scratch, has significantly facilitated the adoption and integration of ML in software systems. Existing ML cloud services cover a wide spectrum of services, from extreme simplicity and complete customizability, for building ML service-based systems (google_cloud_automl, ; ibm_watson, ; ibm_data_platform, ). For instance, some ML cloud services operate as black-box systems, not disclosing the ML classification models they employ, while others offer support throughout ML development stages, from data preprocessing to the selection of ML classifiers, parameter tuning, and model deployment (azure_ml_designer, ). The growing demand for business solutions and the rising popularity of ML services have driven practitioners of diverse skill levels to adopt them to speed up the development, maintenance and evolution of ML-based systems (wan2021machine, ). However, this rapid pace may cause inconsistencies in adhering to best practices for designing and utilizing ML services (wan2021machine, ). As a result, this can lead to service misuse, which can negatively impact the quality of ML service-based systems and pose challenges in their maintenance and evolution (washizaki2022software, ). More specifically, ML service misuse may introduce severe bugs that can affect the correctness, performance, and cost-efficiency of ML service-based systems (wan2021machine, ).
ML service misuse is defined as poor practices in the use of services within the ML development pipeline, including data collection and preprocessing, training, testing, deployment, serving, and monitoring. These misuses are also perceived as a violation of (implicit) service usage constraints in end systems, potentially leading to inefficiencies or unintended consequences on end systems. A recent study reported that 69% of projects on GitHub using Amazon and Google ML cloud services contain service misuses that directly impact the system’s functionality and service costs (wan2021machine, ). For example, failing to specify a stopping criterion during the training of models through an ML service can increase latency as well as cloud usage costs, as training might continue unnecessarily long without significant performance improvements (intro, ). Another example is the availability of synchronous and asynchronous versions of the same ML task offered by several ML cloud providers. Asynchronous services support larger inputs and longer processing times than synchronous services. However, invoking asynchronous ML services in a synchronous manner can drastically impact the performance of an ML service-based system by increasing its latency due to blocked service invocations. Moreover, ML cloud providers typically establish rate limits for ML services, setting the maximum permissible requests within a certain time frame, which helps in regulating ML service utilization and prevents system overloads. However, exceeding such limits without proper handling can result in critical issues, such as increased latency and instability in ML service-based systems. Therefore, it is important to understand ML service misuse in software systems to promote better improvement of software quality and maintain higher levels of correctness, performance, and cost-efficiency.
Unlike ML or service antipatterns, which have well-established definitions in prior research (washizaki2022software, ; bogner2021characterizing, ), ML service misuses still lack precise specifications and a comprehensive understanding in practical contexts. Additionally, existing research on ML service misuses is limited, with the highest number of misuses reported by a single study being eight. (wan2021machine, ). Also, the terminology used by prior research is often inconsistent and have vague definitions, which impedes practitioners from fully understanding and addressing any relevant issues. Recognizing these misuses can help increase awareness of common pitfalls when developing ML service-based systems, thereby potentially improving their overall quality.
Therefore, in this paper, we conduct a comprehensive, multi-vocal empirical study on common misuses in ML cloud services. We propose a catalog of 20 ML service misuses, drawing from a systematic literature review, a gray literature review, and an analysis of ML service-based systems on GitHub. To validate our catalog, we conducted an online survey with 50 ML practitioners, in which we aimed to not only confirm the identified misuses but also to explore their prevalence in industry, understand the underlying causes, and gain insights from practitioners regarding potential mitigation strategies. Our empirical analyses revealed a strong agreement among practitioners on the validity and relevance of our catalog. The results show that ML service misuses are widely spread in industry, with “Improper handling of ML API limits” being the most frequent misuse. Practitioners indicated that these misuses typically arise from a lack of understanding of ML service capabilities, insufficient cloud documentation, and limited awareness of best practices for ML services. To address these challenges, practitioners emphasized the importance of regular performance reviews to maintain efficient ML service-based systems. They also advocated for thorough validation of data quality, combined with continuous education on best practices, to ensure the correct use of ML services.
The contributions of this paper are summarized as follows.
  • We conducted a systematic review of the research literature related to ML systems, cloud services, and (anti)patterns, which resulted in 39 relevant papers, revealing a limited number of ML service misuse scenarios.
  • We conducted a gray literature review of the official documentation provided by major ML cloud providers, which helped identify a more comprehensive set of ML service misuses that have not been studied before.
  • We performed an in-depth empirical analysis of a curated set of 87 GitHub repositories and employed a mixture of manual and static code analysis to identify ML service misuses, resulting in a validated dataset of such misuses.
  • We developed a comprehensive catalog of 20 ML service misuses, categorized by the ML development pipeline stage. This catalog is the largest of its kind and provides a comprehensive overview of potential issues.
  • We conducted a survey with 50 practitioners experienced in ML cloud services, who validated the catalog and provided insights into the use of ML services in the industry. The survey revealed the prevalence of ML service misuses, highlighted their potential causes, and their possible mitigation solution.
The rest of this paper is structured as follows. Section 2 presents the related work. Section 3 describes the study design. The catalog proposal and the analysis of survey results are presented in Section 4. Section 5 discuss the implications of our findings. Section 6 discusses the potential validity threats to our findings. Finally, Section 7 concludes our work and suggests possible future research.
2. Related Work
The use and misuse of ML cloud services remain largely unexplored, with a few exceptions. Wan et al. (wan2021machine, ) analyzed 360 open-source software systems that use Google and AWS ML APIs and identified eight ML API misuses that negatively impact the quality of these systems. However, the authors only analyzed ML cloud APIs that give access to pre-trained models. Unlike our work, ML services misuse related to data preprocessing, model training, and model deployment were not covered. In addition, some misuses were generic in the sense that they apply to any cloud API, not exclusively those related to ML. As a result, in our paper, we considered only four misuses from their study. Moreover, Wei et al. (wei2024demystifying, ) conducted a study on Deep Learning (DL) API misuse in Python libraries such as TensorFlow and PyTorch, exploring their characteristics and proposing an approach to detect them. However, their work was not specific to ML service misuses.
Other research studies addressed the specification of ML (anti)patterns (masuda2018survey, ; washizaki2019studying, ; bogner2021characterizing, ). In particular, Washizaki et al. (washizaki2019studying, ) conducted a literature review and a survey with developers to collect, classify, and discuss (anti)patterns for ML-based systems. The authors found that developers have little knowledge of ML design patterns that could aid in developing their software systems. Bogner et al. (bogner2021characterizing, ) conducted a literature review to study technical debts in ML-based systems and antipatterns (cabral2024investigating, ). The authors argued that there is still no comprehensive conceptual overview of technical debts in ML-based systems, despite existing research efforts. They compiled a set of 72 ML antipatterns aimed at avoiding technical debt in the literature. However, they did not specifically study ML service antipatterns.
With respect to the software quality of ML-based systems, there has been little research in the literature that highlighted the need for adapted quality models (wang2024quality, ; masuda2018survey, ; lewis2021software, ; kuwajima2020engineering, ). For example, Masuda et al. (masuda2018survey, ) summarized software quality methods and techniques for ML-based systems. Kuwajima et al. (kuwajima2020engineering, ) studied the suitability of a traditional software quality model for ML systems and provided insights into the major differences between ML systems and conventional software systems. Cardozo et al. (cardozo2023prevalence, ) analyzed the code quality of reinforcement learning projects by utilizing software metrics related to the definition, access, and interaction of program entities. These metrics served as a proxy of code quality, aiding in the detection of code smells. We leveraged the findings in these studies to specifically target ML service misuses that can negatively impact the quality of software systems. Identifying such misuses would greatly help increase the awareness of common pitfalls when building ML service-based systems and improve their overall quality.
3. Study Design
Figure 1 depicts our study design to build our catalog of ML service misuses. It consists of four steps, which we will describe in details in the following section.
Figure 1. Overview of the Study Design
3.1. Systematic Literature Review
Following the guidelines of Kitchenham et al. (kitchenham2004procedures, ) to conduct systematic literature reviews, we first collected research papers based on a search query. First, we identified key terms related to ML systems, then identified a corresponding set of synonyms that are connected. We thus defined the following search query:
(Machine Learning OR ML OR Artificial Intelligence OR AI OR Deep Learning OR DL) AND
(Misuse OR Antipattern OR Bad Practice OR Pitfall OR Best Practice OR Pattern OR Software Quality OR
Quality Assurance OR Performance) AND (Cloud OR Azure OR Google OR AWS OR Service OR API OR SDK)
Second, we performed this search query on popular digital libraries for research papers, including ACM, IEEE Xplore, and Google Scholar, leading to an initial list of 9,054 papers. The search has been refined by applying filters based on relevance and a publication date range limited to the past ten years. We then filtered these collected papers according to their titles and abstracts. Two of the coauthors manually and independently analyzed all the papers and then reconciled any differences through discussion meetings. We included in our review the papers that met at least two of the following criteria: (1) papers related to ML; (2) papers related to cloud computing services; (3) papers related to (anti)patterns. Finally, we applied forward and backward snowballing techniques (wohlin2014guidelines, ) to minimize the risk of missing relevant papers. Our final set comprises a collection of 39 research papers published between 2017 and 2024. The distribution of these papers across years is depicted in Figure 2. Notably, the highest number of ML service misuses reported in a single study within the existing literature is eight, as provided by Wan et al. (wan2021machine, ). This highlights the need for a more comprehensive and larger catalog of ML misuses, addressing the identified research gap in prior work, as we propose in our study.
Figure 2. Number of Research Papers Published Per Year
3.2. Gray Literature Review
In addition to our systematic literature review, we conducted an in-depth examination of the official documentation provided by major ML cloud providers, including Microsoft, Amazon, and Google. We compiled a detailed list of the ML services offered by each provider and carefully analyzed their documentation. Our focus was twofold: identifying recommended best practices and uncovering non-recommended practices that could affect the performance of ML service-based systems in terms of model accuracy, cloud resource usage, and latency.
To gain deeper insights into the challenges and misuses developers encounter when working with ML services, we also reviewed a wide range of online resources. These included online blogs from the Microsoft Tech Community  [^1] https://techcommunity.microsoft.com, Google Developers [^2] https://developers.googleblog.com/en, and AWS Developer platforms  [^3] https://aws.amazon.com/blogs/developer, as well as publications on Medium and discussions on Stack Overflow. We selected these platforms due to their widespread popularity and extensive repositories of blogs and discussions that frequently address programming challenges in general, and ML-related issues in particular.
We should note that the systematic literature review and the gray literature review provided valuable insights into the practical use of ML services within software systems. By integrating these data, we were able to derive a first set of ML service misuses, which formed a preliminary catalog of 37 misuses.
3.3. Review of Open-Source ML Cloud Service-Based Systems
To collect concrete examples of ML service misuses, we conducted an in-depth analysis of various GitHub repositories. We first started with the dataset provided by Wan et al. (wan2021machine, ) and observed that it contains 307 Python projects covering eight ML service misuses. Moreover, their dataset focused exclusively on ML APIs from Google and Amazon, which limited the scope of ML applications. Given the unavailability of datasets that meet our needs, we proceeded to collect additional GitHub projects, as described below.
  1. a)
    Projects Collection: We collected ML service-based projects on GitHub, a rich source of open-source ML-based software systems. The data collection and analysis was from May 2024 to July 2024, without a specific time frame for the creation date of projects. We used the GitHub API  [^4] https://docs.github.com/en/rest to automatically retrieve projects, enabling a more efficient and structured retrieval of information. We used search keywords listed by cloud service providers. In addition, we employed GitHub’s search functionality to identify the presence of those keywords within Python files. For example, to collect repositories using Microsoft Azure ML, we considered the following keywords: “ML cloud”, “Azure cognitive service”, “API”, “Azure AI”, and “AutoML”. The collected information about the projects includes the repositories’ URLs, number of stars, number of forks, contributors, and descriptions.
  2. b)
    Projects Filtration: We filtered the collected repositories by automatically analyzing the project description and source code. This involved inspecting the development language, identifying the presence of Jupyter notebooks, and locating Python files. This allowed us to filter repositories based on their relevance to our study. The data collection process resulted in 1,900 Python GitHub repositories related to ML services. We randomly selected a sample of these projects (377) for further manual analyses, as we aimed to cover repositories related to major cloud providers (e.g., Microsoft, Google, and Amazon) and ML services (e.g., text-to-speech, and image recognition) to ensure a broad representation. However, we found that not all repositories mentioning cloud services or ML-related terms really use ML services. To maintain the focus on projects that use cloud-based ML services, we excluded repositories that met at least one of the following criteria:
    • Toy projects presented as small demonstrations or tutorials.
    • Official Toolkits from Cloud Providers: These repositories often serve as templates or tools rather than complete projects with fully implemented use cases.
    • Misleading Mention: Projects that mention cloud providers or ML-related terms but do not concretely use any ML service, such as Vertex AI, AWS Sagemaker, or Azure ML, in their source code. For example, we excluded a repository deploying a microservice on Google Cloud but not involving any ML service, despite mentioning Google AI in the ReadMe file.
    Applying the exclusion criteria was done manually and resulted in a curated set of 87 ML service-based projects for further analysis.
  3. c)
    Projects Analyses: Subsequently, two of the coauthors (with experience in machine learning and software engineering) played the role of evaluators and manually and independently analyzed the 87 projects to detect the presence of different ML service misuses. We used Understand [^5] https://scitools.com, a static code analysis tool, to perform a thorough code review and control flow graph generation in complex repositories containing numerous directories. This tool provides different views of the architecture, which enabled a comprehensive understanding and analysis of the system.
  4. d)
    Misuses Isolation: As part of the detection process, the two coauthors independently and systematically identified and highlighted instances of ML service misuse, categorizing them by their stage of ML development pipeline. Each detected misuse is accompanied by a detailed explanation, which provides reasons for why the identified instance is problematic and a link to the exact line of code within the repository, facilitating quick access for further inspection. For example, to detect “Non specification of early stopping criteria” misuse, the two coauthors independently analyzed the ML service-based system, identified the ML service cloud provider used, verified the training component of the code, and reviewed the recommendations for specifying early stopping criteria based on the official documentation of the cloud provider. If the training component did not adhere to best practices for cost-efficiency and resource optimization, a misuse was detected.
  5. e)
    Detection Agreement: The two coauthors tried to resolve any disagreement during the process through discussion meetings. In cases where a consensus could not be reached, a third coauthor was involved to provide additional insights and resolve disagreements. To assess the level of agreement between the evaluators, we calculated Cohen’s kappa coefficient (banerjee1999beyond, ), a statistic used to measure inter-rater reliability. We achieved Cohen’s kappa of 84.7%, indicating a high level of agreement among the evaluators and highlighting the reliability of the manual identification process of the occurrences of ML service misuses.
  6. f)
    Detection Results: Our manual evaluation of the projects on GitHub resulted in the identification of 20 misuses (with 290 occurrences) of ML service misuses in our dataset. To the best of our knowledge, this represents the largest catalog of ML service-based misuses based on various source: systematic and gray literature as well as open-source project analysis (involving both automated and manual). This sets a new benchmark for research in this area. Figure 3 illustrates the prevalence of these misuses, along with the distribution of their occurrences.
Figure 3. Occurrences of ML service misuses in our studied projects, including the misuses identified in the systematic literature [SL] and the gray literature [GL]
3.4. Pre-validation of the Catalog of ML Service Misuses
We curated a preliminary catalog of a total of 34 ML service misuses: 14 misuses were identified through our systematic review of the research literature and 25 through our examination of the gray literature, with an overlap of five misuses. Then, our analysis of GitHub projects uncovered 20 misuses, which included six from the research literature and 14 from the gray literature.
The pre-validation of the catalog involved multiple iterations. The two coauthors independently reviewed and evaluated the relevance of the identified misuses. They applied the following inclusion criteria to retain: (1) misuses specific to ML service-based systems, excluding those applicable to general service systems; (2) misuses applicable to all cloud providers; (3) misuses with occurrences identified within GitHub projects; and (4) misuses that are potentially detectable through source code analysis. They also classified the misuses according to their stage within the ML development pipeline. After such validation, we retained 20 ML service misuses, which we will describe in detail in Section 4.1.
3.5. Survey
We conducted an online survey with practitioners to validate our catalog and get better insight into the usage of ML cloud services in industry. The survey consisted of two main phases:
Preparation of the Online Survey: We created a web-based survey (googleform2025, ) using Google Forms. We designed our survey questions based on the systematic literature review, the gray literature review, and the analysis of ML service-based systems. This helped us identify and shape the different parts of the survey, their various questions, and the response options for each question. Before publishing the survey, we conducted a pilot interview with three professionals and four researchers to identify any duplicated or inconsistent questions in the survey before distribution. This also helped to assess the survey length, which might discourage participants from responding, and to estimate the approximate time required to complete the survey. The seven participants went through the survey and suggested minor modifications. To facilitate participants’ understanding of ML service misuses, we included a link to our online catalog [^6] https://ml-service-misuses.github.io/Catalog-Proposal in the survey to further explain each misuse. The final survey consisted of seven sections, primarily about the demographics of the participants, the frequency of usage of ML services, the introduction of ML misuses, the agreement level with the catalog, and the possible reasons for misuses.
Selection of Participants: We targeted practitioners with experience in ML cloud services. Identifying and inviting such practitioners was challenging due to the key requirement for familiarity with both ML and cloud computing, as well as practical experience with ML services in their work. To select the participants, we relied on (1) information about companies that offer ML cloud services and their clients/partners, (2) the authors’ networks (e.g., former colleagues), and (3) search queries on LinkedIn profiles (e.g., AI Engineer OR Cloud Architect OR Azure certified OR Cloud for AI). For those who were identified as potential practitioners, we sent them invitations via E-Mail, LinkedIn, and Facebook. We chose to limit our invitations to a maximum of five practitioners per organization to ensure a wide representation and mitigate the risk of bias. We invited 421 practitioners to complete the survey and encouraged them to share our invitations within their network. The survey was completed by 50 practitioners, which corresponds to a completion rate of 12%.
Statistical Correlation: We measured the Spearman’s correlation (ali2022spearman, ) between (a) participants’ experience levels in AI development and the number of ML service misuses they agreed upon, (b) participants’ experience levels and the frequency of encountering each ML service misuse. We used the Spearman’s correlation as it does not require assumptions about data distribution.
4. Results
4.1. Catalog of ML Cloud Services Misuses
We categorized the ML service misuses according to the stages of the ML development pipeline and published our catalog proposal online (catalog, ). Below, we provide a detailed description of each misuse, along with illustrative examples of some misuses derived from our analysis of GitHub repositories.
  1. 1)
    Data Collection & Preprocessing Stage
    Misuse 1 - Not using batch API for data processing: Cloud providers offer batch processing APIs to optimize data loading performance by handling data in batches. However, developers sometimes fail to use these batch APIs, opting instead to load data items individually or implement their own batch processing solutions (cao2022understanding, ). Not using a batch API can cause out-of-memory issues, excessive network traffic, and significant delays in data loading, ultimately slowing down model training and increasing operational costs. The code below  [^7] https://github.com/ovokpus/Python-Azure-AI-REST-APIs/blob/main/text-analytics-sdk/text-analysis.py shows two occurrences of this misuse, as the Azure Text Analytics batch API was not used for language detection and sentiment analysis across multiple documents (Batch, ), leading to slower data processing. Below, we present a refactored version that uses batch processing, reducing redundant calls and network traffic.
    Misuse 1: Not using Batch API for Data Processing (Example)
    from azure.ai.textanalytics import TextAnalyticsClient
    def main():
    cog_client = TextAnalyticsClient(endpoint=cog_endpoint, credential=credential)
    # Analyze each text file in the reviews folder
    for file_name in os.listdir(reviews_folder):
    # Read the file contents
    detectedLanguage = cog_client.detect_language(documents=[text])[0]
    sentimentAnalysis = cog_client.analyze_sentiment(documents=[text])[0]
[…]
    Misuse 1: Not using Batch API for Data Processing (Recommendation)
    from azure.ai.textanalytics import TextAnalyticsClient
    def main():
    # Load documents directly from files for batch calls
    documents = [open(os.path.join(reviews_folder, file_name), encoding=’utf8’).read() for file_name in os.listdir(reviews_folder)]
    # Batch process for language and sentiment detection
    detected_languages = cog_client.detect_language(documents=documents)
    sentiments = cog_client.analyze_sentiment(documents=documents)
    Misuse 2 - Inefficient data transmission: Refers to inefficient data transmission between components within an ML service-based system, such as storage services, compute nodes, and other cloud resources. This inefficiency leads to increased latency and higher cloud usage costs (cao2022understanding, ). For example, the repeated transfer of training data from cloud storage to compute nodes for each job, instead of caching it locally, can cause significant delays in training time, increased network traffic, and higher data transfer costs.
  2. 2)
    Training Stage
    Misuse 3 - Non specification of early stopping criteria: ML services typically offer options to set early stopping criteria, helping to avoid overfitting and unnecessary computational expenses (microsoft2024hyperparameters, ). However, developers sometimes fail to specify these criteria, which allows the training to proceed for more epochs than necessary. This oversight can result in wasted computational resources, increased training duration, higher costs, and potential overfitting (ying2019overview, ).
    Misuse 4 - Avoiding parallel training experiments: Cloud providers offer the capability to run parallel training experiments to accelerate the model training process and enhance the efficiency of ML service-based systems (microsoft_parallel_models, ). Avoiding parallel experiments can slow down model development and prevent the full utilization of available computational resources. In the following example [^8] https://github.com/praneet22/DevOpsForAI/blob/master/aml_service/10-TrainOnLocal.py, the Azure ScriptRunConfig class is used without specifying a distribution configuration, which effectively disables parallel training, resulting in a less efficient process. To improve efficiency, it is recommended to either (1) include a distribution configuration within ScriptRunConfig (as shown in the listing below) or (2) use Azure DistributedRunConfig, which automatically distributes computation across multiple nodes, thereby enabling parallel training experiments and significantly accelerating the training process (scriptrunconfig, ).
    Misuse 4: Avoiding Parallel Training Experiments (Example)
    from azureml.core import Experiment
    from azureml.core import ScriptRunConfig
    src = ScriptRunConfig(
    source_directory="./code",
    script="training/train.py",
    run_config=run_config_user_managed,)
    run = exp.submit(src)
    Misuse 4: Avoiding Parallel Training Experiments (Recommendation)
    from azureml.core import Experiment
    from azureml.core import ScriptRunConfig
    # Create a distributed run configuration
    distributed_run_config = DistributedRunConfig(
    node_count=2, # Number of nodes
    virtual_environment=’env1’,
    instance_type=’Standard_NC6’, #VM size)
    src = ScriptRunConfig(source_directory="./code",
    script="training/train.py",
    run_config=run_config_user_managed,
    distributed_job_config=distributed_run_config)
    Misuse 5 - Not using automatic hyperparameter tuning: ML cloud providers offer the capability to define the search space and automatically optimize ML model hyperparameters (microsoft2024hyperparameters, ). However, developers might not leverage this feature, choosing instead to manually set hyperparameters, which can lead to suboptimal model performance and increased training time (victoria2021automatic, ; weerts2020importance, ).
    Misuse 6 - Not using training checkpoints: Cloud providers offer the ability to resume training from the most recent checkpoint, saving the current state of the experiment rather than starting from scratch. This can save significant time and computational resources, especially when training large and complex models (xu2022checkpointing, ; rojas2020study, ). However, developers may neglect to save training checkpoints in cloud storage. If a model fails and checkpoints have not been saved, the entire training job or pipeline will terminate, resulting in a loss of data since the model’s state is not preserved in memory (google2024mlbestpractices, ). In the following example [^9] https://github.com/Subrahmanyajoshi/Cancer-Detection-using-GCP/blob/master/detectors/detector.py, the training process in Google Vertex AI commences without implementing checkpoints to save the model’s state. To prevent potential data loss and improve efficiency, a save_checkpoint function can be defined (as shown in the recommendation listing) to store the model’s state in Google Cloud Storage during training.
    Misuse 6: Not Using Training Checkpoints (Example)
    from detectors.vertex_ai_job import Trainer
    # Run Trainer
    if args.train:
    print(’Initialising training’)
    trainer = Trainer(config=config)
    trainer.train()
    Trainer.clean_up()
    Misuse 6: Not Using Training Checkpoints (Recommendation)
    from detectors.vertex_ai_job import Trainer
    from google.cloud import storage
    #Save checkpoints in Google Cloud storage
    def save_checkpoint(checkpoint_path)
    storage_client = storage.Client()
    # Define a Google Cloud Storage container
    bucket = storage_client.bucket("MyBucket")
    # Create a blob from the bucket
    blob = bucket.blob(checkpoint_path.split("/"))
    # Upload the file to Google Cloud Storage
    blob.upload_from_filename(checkpoint_path)
    print(f"Checkpoint saved to {blob.name}")
    # Run Trainer
    if args.train:
    print(’Initialising training’)
    trainer = Trainer(config=config)
    trainer.train()
    trainer.save_checkpoint("path/to/save")
    Trainer.clean_up()
    Misuse 7 - Bad choice of training compute targets: Refers to selecting non-optimal hardware and compute resources for training ML models. Cloud providers offer various types of training compute targets, yet not all resources are adequate for every ML task. For example, Azure Machine Learning Kubernetes is ideal for ML pipelines and Azure ML Designer, which require robust support for parallel processing and resource allocation. However, it is not recommended for automated ML, which requires different computing capabilities (azure_compute_targets, ). Choosing the right compute target ensures optimal performance and cost-effectiveness for the specific training task.
    Misuse 8 - Excluding algorithms in automated ML: Cloud providers offer automated machine learning services that perform experiments with various ML algorithms to generate an optimized model for a given prediction task ready for deployment (microsoft2024autotrain, ). However, when configuring these services, developers may mistakenly exclude promising algorithms, thereby limiting the model’s effectiveness and performance.
  3. 3)
    Testing Stage
    Misuse 9 - Misinterpreting output: ML cloud services offer pre-built models that operate on high-dimensional continuous representations but often produce a small discrete set of outputs. Consequently, the output of these services may contain complex and easily misinterpretable semantics, potentially leading to bugs (wan2021machine, ). For example, Google’s NLP Sentiment Detection API returns two metrics: score and magnitude, which together assess the sentiment of a text input. Developers may misinterpret the sentiment if they do not correctly combine both metrics. The score reflects the polarity of the sentiment (positive or negative), while the magnitude indicates the strength of the sentiment (wan2021machine, ). The code example below, from a journal application [^10] https://github.com/beekarthik-ai/JournalBot/blob/master/webapp/app_folder/routes.py, judges the sentiment in a user’s journal. However, it only considers the score value, potentially marking the journal as emotionally negative incorrectly. As mentioned in Google’s service documentation, it is recommended to extend the sentiment assessment with a rule-based evaluation using both score and magnitude to predict the final sentiment, as shown below.
    Misuse 9: Misinterpreting Output (Example)
    from google.cloud import language_v1
    response=client.analyze_sentiment(document^document,
    encoding_type=encoding_type)
    sentiment=response.document_sentiment.score
    if avg_sentiment < 0:
    message= ’Your posts show that you might not be
    going through the best of time.’
    elif avg_sentiment<0.5:
    message = ’Your posts show that you are having some decent times! Here s to many more happy days’
    Misuse 9: Misinterpreting Output (Recommendation)
    from google.cloud import language_v1
    # Assuming ‘client‘, ‘document‘, and ‘encoding_type‘ are already defined
    response = client.analyze_sentiment(document=document, encoding_type=encoding_type)
    score = response.document_sentiment.score
    magnitude = response.document_sentiment.magnitude
    if abs(score) < 0.15 or abs(magnitude) < 0.15:
    message = "Neutral"
    elif score > 0:
    message = "Here’s to many more happy days"
    else:
    message = "You might not be going through the best of times."
    print(message)
    Misuse 10 - Ignoring fairness evaluation: ML cloud providers offer mechanisms for fairness evaluation, which is important to ensure unbiased and equitable models (mehrabi2021survey, ; antunes2018fairness, ). Neglecting fairness evaluation can result in predictions that have potential biases and unfair towards specific demographic groups (aws2024mllens, ), which could negatively impact model’s generalizability and performance.
    Misuse 11 - Ignoring testing schema mismatch: Cloud providers offer ML services to detect unmatched data schemas, which include feature or data distribution mismatches between training, testing, and production data, often by raising alerts. However, developers may ignore setting up these alerts or disable them. For example, Amazon ML displays alerts if the schemas for the training and evaluation data sources are not consistent (aws2024mlevaluationalerts, ). Disabling these alerts can result in missing discrepancies, such as features present in the training data but absent in the evaluation data, or detecting unexpected additional features. This oversight may weaken the model’s accuracy and performance in the production environment (polyzotis2019data, ).
    Misuse 12 - Using suboptimal evaluation metrics: Some ML services optimize and evaluate models based on specified evaluation metrics. These metrics determine how model performance is measured during training and evaluated during testing (naser2020insights, ). However, developers sometimes choose suboptimal evaluation metrics, leading to less effective models that do not align well with business needs or dataset characteristics. For example, according to Microsoft ML’s official documentation (microsoft2024autotrainconfig, ), in classification tasks, metrics such as accuracy and norm_macro_recall may not perform well with small datasets, datasets with significant class imbalance, or when metric values are close to zero or one. In such cases, AUC_weighted could be a better option (microsoft2024autotrainconfig, ).
  4. 4)
    Deployment Stage
    Misuse 13 - Overwriting existing ML APIs without versioning: Cloud providers offer ML API versioning through tools such as Azure API Management and AWS Management Console. Without version control, it becomes challenging to track changes, revert to previous versions, or understand the evolution of the deployed model. Developers may, however, overlook these tools and unintentionally overwrite existing ML APIs without proper versioning, which can lead to issues in the production environment. As discussed in a Stack Overflow post (stackoverflow_ci_cd_ml_studio, ), when publishing an updated model using Azure ML Studio, overwriting an existing API without versioning can make it difficult to roll back to a previous stable version if the new model introduces unexpected behavior or lower accuracy, potentially causing disruptions in the production system.
    Misuse 14 - Choosing the wrong deployment endpoint: Cloud providers offer online and batch endpoints for deployment. Online endpoints are mainly used to operationalize models for real-time inference in synchronous, low-latency requests. In contrast, batch endpoints are used primarily to operationalize models or pipelines for long-term, asynchronous inference (MicrosoftAzureMLEndpoints, ). However, developers may select the inappropriate endpoint for deployment. For example, using a batch endpoint for real-time tasks can delay inferences since batch processing is not optimized for speed.
    Misuse 15 - Disabling automatic rollbacks: ML cloud service providers offer features that automatically rollback to a previously stable version of a model if errors or performance issues arise with a newly deployed version. However, developers may disable this feature, allowing poorly performing models to remain in production, which can negatively impact system performance (serban2020adoption, ).
    Misuse 16 - Disabling automatic scaling for online prediction service: Automatic scaling allows resources to be dynamically adjusted based on demand to maintain sufficient capacity for online prediction services. This feature helps manage varying prediction request rates while minimizing cloud usage costs (google2024scalingmlpredictions, ; google2024mlgcpbestpractices, ). However, developers may disable automatic scaling when deploying ML models, potentially increasing inference latency during peak demand periods. The example below  [^11] https://github.com/Ewen2015/BatCat/blob/master/batcat/SageMaker.py highlights this issue since automatic scaling is disabled by default in Amazon SageMaker. As a result, developers need to explicitly configure auto-scaling policies. It is recommended to use SageMaker’s “AutoScalingPolicy” to automatically adjust resources according to demand, ensuring optimal performance during traffic spikes and minimizing costs during low usage periods (as shown in the recommendation snippet) (appawsautoscaling, ; onlineendpoints, ).
    Misuse 16: Disabling automatic scaling for online prediction service (Example)
    import sagemaker
    endpoint_config_response = client.create_endpoint_config(
    EndpointConfigName=epc_name,
    ProductionVariants=[ {"VariantName": "sklearnvariant",
    "ModelName": model_name,
    "InstanceType": "ml.c5.large",
    "InitialInstanceCount": 1 },],)
    Misuse 16: Disabling automatic scaling for online prediction service (Recommendation)
    import sagemaker
    endpoint_config_response = client.create_endpoint_config(
    EndpointConfigName=epc_name,
    ProductionVariants = [{"VariantName": "sklearnvariant",
    "ModelName": model_name,
    "InstanceType": "ml.c5.large",
    "InitialInstanceCount": 1,
    "AutoScalingPolicy": {
    "ScaleInCooldown": 60, # waiting time before next scale
    "MinCount": 1, # Minimum number of instances
    "MaxCount": 3 }}]
  5. 5)
    Serving Stage
    Misuse 17 - Improper handling of ML API limits: Failure to adhere to API request rate limits can compromise the stability and performance of the ML service (MicrosoftTechCommunity2024, ). Developers may not adequately manage these limits, causing predictions to abruptly halt when the rate is exceeded. For instance, surpassing the maximum number of API calls within a set timeframe, such as requests per second defined by the Azure OpenAI service (MicrosoftOpenAIQuotasLimits, ), can result in delayed or rejected requests until they conform to the permitted rate.
    Misuse 18 - Misusing Synchronous/Asynchronous ML APIs: ML cloud providers typically offer both synchronous and asynchronous versions of the same tasks. Asynchronous services are designed to handle larger inputs and longer processing times, as opposed to synchronous services. However, calling asynchronous ML services in a synchronous manner can degrade system performance significantly, as it increases latency due to blocked service invocations. (wan2021machine, ).
    Misuse 19 - Calling the wrong ML service API: Cloud providers often offer multiple ML APIs for similar tasks. Without a thorough understanding of these APIs, developers might use the wrong one, leading to significantly reduced prediction accuracy, incorrect prediction results, or software failures. For example, image classification and object detection are vision APIs that provide description tags for input images. The former offers a single tag for the entire image, while the latter provides a tag for each object. Using image classification instead of object detection can cause software to miss important objects, and using object detection instead of image classification can result in incorrect image tags (wan2021machine, ).
  6. 6)
    Monitoring Stage
    Misuse 20 - Ignoring monitoring for data drift: This refers to neglecting the continuous assessment of changes in statistical characteristics or data distributions, which is crucial for maintaining model performance (bova2017intervention, ). Data drift occurs when the incoming data distribution differs from the training data, leading to degraded model accuracy over time (mallick2022matchmaker, ). Cloud providers recommend implementing skew and drift detection mechanisms to monitor these changes and alert developers when significant changes occur (google2024mlgcpbestpractices2, ). By detecting data drift early, models can be retrained or adjusted to ensure they continue performing as expected in production environments.
4.2. Analysis of the Survey Results
In this section, we present the survey results, with data collected from responses between July and December 2024.
Survey Participants: We reached a diverse group of 50 practitioners who actively participated in ML service-based projects, representing a wide range of roles within the field. Specifically, as shown in Figure 4, 34% of the participants identified as ML engineers, another 34% as data scientists, and 16% as software engineers. The remaining 16% of the participants held various other roles, including cloud engineers, data engineers, and software architects, highlighting the breadth of expertise and perspectives involved in the survey. In terms of experience, our participants also reflected considerable diversity. The majority reported having between one to five years of experience in AI (56%), software development (52%), and cloud computing (42%). Furthermore, 16% of the practitioners had less than one year of experience in both AI and software development, and 42% had less than one year of experience in cloud computing. This variety in experience levels further underscores the diversity of the participants, encompassing both emerging professionals and those with more established expertise in ML service-based systems.
Figure 4. Roles of Survey Participants
[…]
Figure 5. Projects Distribution
Figure 6. Projects Scale
We asked practitioners about the frequency of using ML services. Figure 7 indicates that these services are commonly used in the industry, with 52% reporting monthly to daily usage, whereas only 16% use them once a year or less. This frequent use highlights the necessity for best practices. We also asked about which services are used most often and found Azure ML services, Amazon ML services, and Google ML services are at the top. Surprisingly, only 10% of practitioners use IBM Watson, and 6% use services from Oracle or others. Practitioners interact with ML services using various methods: 70% use APIs, 66% use Command Lines (CLI) and User Interfaces (UI), and 62% use SDKs, as shown in Figure 8. A preference for APIs is noted due to their integration ease, while the high usage of CLIs and UIs suggests a balance between technical and non-technical engagement.
We should note that diversity in (1) project sizes based on ML services, (2) the degree of ML services adoption, (3) interaction methods, and (4) different cloud providers enhances the generalizability of our findings, despite the number of participants in our survey.
Figure 7. Frequency of ML Service Utilization
Figure 8. Methods of ML Service Utilization
ML Services Impact and Challenges: Practitioners highlighted both positive and negative impacts of ML services on their project development process. On the positive side, 82% of practitioners reported that adopting ML services in their industrial projects helped them save development time. Also, 68% found that their projects were more scalable when using ML services, and 56% noted an increase in the efficiency of the software development process, as remarked in Figure9. These findings underscore the critical role of ML services in enhancing system scalability and efficiency while reducing the time required for development.
Practitioners have identified several challenges associated with ML services. Over half (54%) reported integration difficulties, indicating ML services can complicate the incorporation of new components into existing systems. Resource management issues were highlighted by 46% of practitioners, who experienced problems with memory and computational resource allocation. Security concerns were also prevalent, with 42% of participants expressing worries about data security and privacy. These challenges emphasize the need for effective practices to mitigate the negative impacts of ML services. In addition, 77% of practitioners faced challenges at various stages of their projects. Specific difficulties included data collection and processing (42%), training (38%), testing (35%), deployment (50%), serving (38%), and monitoring (40%). These findings illustrate the widespread nature of the challenges related to ML services and the need for comprehensive strategies to address them.
Figure 9. Impact of ML Services
Figure 10. ML Service-Related Challenges
Figure 11. Frequency of ML service challenges across different stages of the ML development pipeline
Takeaway 1: To fully unlock the benefits offered by ML services, it is important to address the challenges related to integration, resource management, and security. Tackling these issues will not only make projects succeed, but also boost significant benefits like efficiency, scalability, and improved capabilities in project development.
Catalog Agreement Analysis: We asked practitioners to what extent they agree or disagree that the practices in our catalog represent misuses of ML services. Figure 12 illustrates the level of agreement for each identified misuse. We found that all practices were considered misuses, with agreement levels ranging from 52% to 80%. The highest consensus (80%) was on misuses such as “Choosing the wrong deployment endpoint”, “Overwriting existing ML APIs without versioning”, “Ignoring fairness evaluation” and “Misinterpreting output”. On the other hand, the lowest agreement (52%) was for “Not using Batch API for data processing”. This lower score may reflect that some practitioners did not see this as a misuse, particularly in scenarios where small datasets make batch processing unnecessary. Furthermore, some practitioners were neutral, potentially due to varying contexts and use cases influencing their perceptions of these practices as misuses.
Takeaway 2: There is a strong consensus on the significance of our ML service misuse catalog, with the majority of practitioners acknowledging its crucial impact. This highlights a broad recognition of practical challenges within the industry, emphasizing the urgent need for strategic interventions and solutions.
Figure 12. ML Service Misuses Agreement Level
Correlation between Experience Levels and Agreement on ML Service Misuses: Our analysis of the correlation between participants’ experience levels in AI development and the number of ML service misuses they agreed upon revealed a moderate positive and statistically significant correlation (48%) between experience levels and agreement on ML service misuses. In addition, we studied the correlation between the participants’ experience level and their agreement level with each misuse. In 15 out of 20 misuses, we found positive and statistically significant correlations ranging from 30% to 47%. The positive correlation indicates that as developers gain more experience in AI development, they are more likely to agree on potential misuses in ML services, enabling them to identify misuse scenarios more effectively.
ML Service Misuses Frequency: We asked practitioners how frequently they encounter or tend to introduce each of the ML service misuses listed in our catalog. Specifically, for each misuse, practitioners were asked to indicate whether they have encountered the misuse in most ML projects (more than 75% of projects), in some ML projects (less than 75% of projects), or never encountered the misuse. To ensure reliability, for each misuse, we only retained responses from practitioners who only agreed on the misuse. As illustrated in Figure 13, our findings reveal that ML service misuses are prevalent in the industry, with only 10% to 32% of practitioners reporting that they had never encountered the misuses studied in their projects. We also found that “Improper handling of ML API limits” has the highest percentage of being encountered in most projects (31%). This could be due to the poor documentation available for setting proper ML API limits from cloud providers.
In addition, “Inefficient data transmission” has the highest percentage of being encountered in some projects (70%). This may reflect the common challenge of using caching mechanisms, for example, to process redundant data transfers. “Not using Batch API for data processing” is also a common problem in some projects (68%). This reflects the lack of awareness regarding the advantages of batch processing, resulting in the adoption of less optimal data processing methods. On the other end, “Excluding algorithms in automated ML” has the highest percentage of never being encountered (32%), followed by “Calling the wrong ML service API” (26%). The above findings suggest that these specific misuses are less common in practice.
Figure 13. ML Service Misuses Frequency
Takeaway 3: Misuses of ML services are prevalent across various stages of project development, with the most common issue being the improper handling of ML API limits. This oversight can lead to significant performance bottlenecks and service interruptions, thus addressing it becomes crucial to ensure smooth operations.
Correlation between Experience Levels and Frequency of ML Service Misuses: In nine of 20 misuses, we found positive and statistically significant correlations ranging from 29% to 36% between the level of experience of participants and the frequency of encountering each ML service misuse. This suggests that more experienced practitioners are better at identifying and recognizing these issues. This could indicate that more practiced professionals have a deeper understanding of ML service-based systems, are more familiar with common pitfalls, or are exposed to complex projects where such misuses are more prevalent.
Reasons of ML Service Misuses: We asked the practitioners about the possible reasons behind the presence of ML service misuses in practice. We included reasons in the survey based on our literature review (nahar2022collaboration, ; nahar2023meta, ; bogner2021characterizing, ) and provided an “Other” option to allow participants to add additional reasons freely. As shown in Figure 14, the most prominent issues identified were a lack of understanding of ML service capabilities (58%) and lack of documentation (50%). These findings highlight the urgent need for better educational resources and comprehensive documentation from ML cloud service providers. Also, half of the practitioners pointed to a lack of awareness or understanding of best practices for using ML cloud services and poor code review processes (44%) as significant reasons for introducing ML service misuses. This underscores the importance of disseminating best practices and establishing robust code review mechanisms. Other mentioned factors included pressure to meet tight deadlines (34%), and the rapid evolution of ML cloud services, which complicates staying updated with the latest changes (30%). As shown in Figure 14, no “Other” reasons were mentioned, which highlights the wide coverage of the predefined reasons in the survey. These insights suggest that addressing ML service misuses will require a multi-faceted approach, including improved documentation, and better education on ML capabilities, dissemination of best practices, and enhancing resource availability and expertise.
Takeaway 4: ML service misuse mainly stems from a lack of understanding of their capabilities, insufficient documentation, and limited awareness of best practices, which could lead to inefficiencies and poor performance. Hence, it is crucial to foster greater awareness and provide clear guidance to ML practitioners in this context to ensure more reliable and efficient deployments.
Figure 14. Reported Reasons of ML Service Misuses
Mitigation of ML Service Misuses: We asked the practitioners for mitigation solutions to avoid introducing ML service misuses.As shown in Figure 15, practitioners highly recommend to stay updated with best practices of ML services and cloud providers standards, as mentioned by 66% of the survey participants. This ensures that practitioners will be aware of the latest guidelines, tools, and techniques, reducing the likelihood of introducing potential ML service misuses. Furthermore, 64% of practitioners emphasized the importance of ensuring thorough validation of the quality of data used by ML components and conducting regular model and system performance reviews (58%). Other recommended practices include implementing robust alerting systems for model monitoring (48%), effective collaboration with domain experts and stakeholders (48%), and active engagement in code peer reviews (48%). These strategies collectively, reduce the risk of introducing ML service misuses, thus enhancing the quality and reliability of ML service-based systems.
Takeaway 5: To effectively mitigate ML service misuse, ML engineers should: (1) continually align with the latest best practices and standards from ML services and cloud providers, (2) rigorously validate ML data to ensure quality, and (3) systematically perform regular performance evaluations to detect issues early. These steps help maintain reliability and model accuracy.
Figure 15. ML Service Misuses Mitigation
Table 1. Summary of the Developed Catalog
ML Development Misuse Name Example Recommendation Frequency% Agree%
Pipeline Stage Most Some
1 Data Collection Not using Batch API for data processing Processing multiple documents with separate API calls Load data in batches to optimize performance and prevent OOM issues 22% 68% 52%
2 & Preprocessing Inefficient data transmission Repeated data transfers instead of caching Cache data locally to cut transfer time and costs 20% 70% 57%
3 Training Non-specification of early stopping criteria Overtraining without stopping criteria Set early stopping to reduce training time and avoid overfitting 24% 56% 72%
4 Avoiding parallel training experiments Missing distribution configuration Enable parallel training with a distribution config 18% 66% 54%
5 Not using automatic hyperparameter tuning Manually tuning hyperparameters Use automated hyperparameter tuning for efficiency 16% 68% 62%
6 Not using training checkpoints No checkpointing in training Save checkpoints to resume training and prevent data loss 14% 64% 68%
7 Bad choice of training compute targets Using non-recommended Azure ML Kubernetes Choose compute targets suited to the ML task 18% 66% 78%
8 Excluding algorithms in automated ML Ignoring a promising algorithm Include all relevant algorithms for better performance 16% 52% 57%
9 Testing Misinterpreting the output Focusing on ’score’ while ignoring ’magnitude’ Use rule-based evaluations with multiple metrics 20% 62% 80%
10 Ignoring fairness evaluation Not checking demographic biases Ensure fairness with bias evaluation tools 26% 58% 80%
11 Ignoring testing schema mismatch Disabling Amazon ML schema alerts Set alerts for data schema mismatches 18% 64% 74%
12 Using suboptimal evaluation metrics Using accuracy instead of AUC in imbalanced data Select evaluation metrics suited to the task 18% 60% 78%
13 Deployment Overwriting existing ML APIs without versioning No model versioning, preventing rollback Use ML API versioning for tracking and rollback 18% 66% 80%
14 Choosing the wrong deployment endpoint Using batch endpoints for real-time tasks Use online endpoints for real-time and batch for async tasks 20% 64% 80%
15 Disabling automatic rollbacks Disabling automatic deployment rollbacks Enable auto rollbacks for stable production models 20% 60% 72%
16 Disabling auto-scaling for online prediction service No auto-scaling in Amazon Sagemaker Configure auto-scaling for dynamic resource allocation 20% 66% 54%
17 Serving Improper handling of ML API limits Exceeding API call limits Set API rate limits for stable performance 31% 45% 72%
18 Misusing Sync/Async ML APIs Calling asynchronous ML services synchronously Use async APIs for large inputs, sync for low-latency tasks 18% 60% 68%
19 Calling the wrong ML service API Using image classification instead of object detection Choose the right API for the task 14% 60% 76%
20 Monitoring Ignoring monitoring for data drift Not setting up alerts for data drift Detect and address skew and drift for model reliability 26% 58% 72%
5. Discussion and Recommendations
Towards a better understanding of ML service misuses. We present the largest catalog of 20 ML service misuses (summarized in Table 1), which we identified through a multi-vocal study and a survey of 50 practitioners. We identify 290 occurrences of misuses based on the analysis of a curated set of 87 ML service-based systems. Our study does not only map out these misuses in the literature and ML service-based systems, but also quantifies practitioners’ agreement with our catalog and the frequency of encountering these misuses in the industry. We should note that the disparity between frequency (e.g., 68%) and agreement (e.g., 52%) for certain misuses, such as “Not using batch APIs for data loading”, highlights that while frequency indicates how often participants encounter the issue, agreement reflects a shared understanding of its problematic nature. This gap suggests that observing a misuse does not guarantee consensus on its consequences, likely due to variations in participants’ interpretations, project contexts, or professional backgrounds. Also, our survey results indicate that a significant percentage of participants have frequently encountered specific misuse in their professional experience, such as “Improper Handling of ML API limits” (31%). However, our analysis of open-source projects revealed only a few occurrences (30) of this misuse, highlighting a notable discrepancy.
There could be several reasons for this gap. Firstly, survey participants may work with production-level ML systems, where misuses are more common but often remain undocumented in public repositories. In addition, many ML misuses occur in private workflows, deployment pipelines, or closed-source production environments, making them less visible to open-source communities, such as GitHub, which was the basis for our analysis. Future research should explore this disconnect by examining industry case studies or interviewing practitioners to understand how such misuses are managed in real-world applications. Integrating insights from practitioners with open-source analysis would provide a comprehensive view of ML service misuses and their broader implications.
Moreover, the Spearman’s correlation between the frequency of encountering ML service misuses in most projects and their occurrence in GitHub projects is 0.46 (p-value = 0.04), indicating a moderate, positive, statistically significant correlation. This implies that as the frequency of ML service misuses increases, their occurrence across GitHub projects also tends to rise. The statistical significance (p ¡ 0.05) confirms that this relationship is unlikely to be due to random chance. This finding suggests that certain antipatterns are not only widely recognized and analyzed within software development but are also commonly observed in real-world projects. This prevalence can potentially affect the software quality and maintainability of software.
Towards more awareness of ML service misuses. The positive correlation between the experience level and the frequency of encountering those misuses, highlights the importance of knowledge transfer from experienced practitioners to less experienced ones, emphasizing the need to raise awareness about the good practices of ML services usage and improve the overall quality of ML service-based systems. This study has significant implications for various stakeholders in the ML ecosystem, including software/ML engineers, ML cloud providers, and open-source platforms like GitHub.
    For practitioners, our catalog of ML service misuses offers valuable insights into common pitfalls and misconfigurations that can reduce the reliability of ML service-powered systems. With this comprehensive catalog, practitioners can better understand potential risks and proactively design and integrate ML services that align with best practices. For example, software engineers, who often serve as the bridge between model development and production environments, can leverage this catalog to mitigate risks related to fairness, scalability, and performance.
    For cloud ML service providers, our study underscores the importance of developing tools and platforms that facilitate the detection and prevention of ML service misuse. We observed that some ML service misuses are trivial and could easily be spotted by service providers (e.g., “Overwriting existing ML APIs without versioning”). Hence, such tools would help ensure that users of ML services have access to transparent, scalable, and high-performance systems.
    Platforms such as GitHub and other open-source communities can play a critical role in promoting collaboration to identify, address, and mitigate these misuses, driving the creation of shared resources that increase the performance of ML service-based systems. For instance, GitHub could integrate a tool that detects ML service misuses when practitioners submit a pull request or commit their work, helping to flag potential misuses early in the development process.
Overall, our findings trigger an alarm towards investing in more robust tooling, policies, and guidelines to safeguard the integrity and societal impact of ML cloud services in particular and ML based systems in general.
Towards new methodologies for ML service misuses detection. Our study establishes a foundation for developing automated detection and refactoring tools to enhance the quality of ML service-based systems. While some ML misuses can be easy to detect and refactor (e.g., “Non specification of early stopping criteria”, “Misinterpreting output”), others (e.g. “Ignoring fairness evaluation”, “Ignoring monitoring for data drift”) can be more complicated and require dedicated techniques to address. By cataloging prevalent ML service misuses and their occurrences in open-source projects, we provide a concrete basis for building tools that automatically identify and refactor these issues. Future research should focus on extending and refining this catalog, and developing solutions that help practitioners adhere to best practices, thereby improving the quality and maintainability of such systems.
6. Threats To Validity
We describe in this section the threats to the validity of our study.
6.1. Construct validity:
When considering the construct validity of our study, some threats must be acknowledged. One key concern is the method used to collect GitHub projects with identified misuses. Our scripts for project identification may have unintentionally excluded or included certain projects that were not entirely relevant to the scope of our study. To mitigate this risk, we conducted a manual review of the projects to ensure their relevance to our research on ML service misuses. This additional step helped refine the dataset and enhance its accuracy.
Furthermore, while our catalog of 20 misuses is based on a thorough review of both research and gray literature, there is always a possibility that some misuses were overlooked. Despite its comprehensiveness, the catalog may not capture every potential misuse in ML services. For instance, platforms such as Stack Overflow could offer valuable insights into other crucial misuses not yet addressed in our catalog. To address this, we are extending the catalog in a parallel independent study, incorporating contributions from such platforms to further strengthen its completeness and relevance in the field of ML service misuses.
6.2. Internal validity:
Our search query might not cover all terms related to ML service misuses and could miss important research resources. We accept these threats as our goal was not to provide an exhaustive list of all existing ML service misuses. As a mitigation solution, we (1) included in our search query the most important keywords related to the software quality of ML service-based systems, and (2) applied forward and backward snowballing to minimize the risk of missing important resources.
Furthermore, social desirability is a bias that leads our survey participants to agree with misuses. To minimize this threat, we did not offer any incentives for practitioners to participate in our survey. We also guaranteed the practitioners their anonymity and emphasized that all the reported information would be for research purposes only.
Another internal threat to the validity of our study is the practitioners’ interpretation of the identified misuses. To mitigate this, we provided a link to our online catalog within the survey, allowing participants to access detailed descriptions of each misuse. In addition, we categorized the misuses according to their respective stages within the ML development pipeline, ensuring a structured understanding. The catalog also includes real-world examples of misuse occurrences, along with proper references, to facilitate a more informed and accurate assessment of the misuses. We also extended invitations for interviews to practitioners to collect additional insights. However, we did not receive any responses.
One potential threat to internal validity is the manual analysis, which could introduce bias in the findings. To mitigate this, two evaluators (coauthors of this paper, with experience in machine learning and software engineering) independently analyzed the projects based on the same criteria. All discrepancies were resolved through discussion and consensus meetings. The agreement level between the evaluators was measured using Cohen’s Kappa and showed a strong level of consistency, reducing the risk of bias in the analysis.
6.3. External validity:
These threats addresses the generalizability of our findings. While we surveyed 50 practitioners, we highly believe that this sample size is reasonable considering the length of the survey and the challenge of identifying professionals with specific experience in ML cloud services. Importantly, the participants represent a diverse group of practitioners, all actively involved in ML service-based projects. This diversity, coupled with the hands-on experience of the participants, strengthens the reliability of our findings as it reflects the insights of subject matter experts.
Our study focuses on a sample of a curated set of 377 GitHub projects, all written in Python. While this sample provides valuable insights into ML service misuses, it may not fully represent the broader landscape of ML service misuses. The limited scope in terms of both project numbers and programming language means that the findings might not be directly generalizable to other programming languages. Future research should on expanding this sample to include a more diverse range of projects across different programming languages to enhance the generalizability and applicability of the results.
Another external threat pertains to the coverage of reasons and mitigation solutions for ML misuses. To mitigate this threat, we mentioned possible reasons and mitigation solutions in the survey based on our literature review and practical experience. Before disseminating the survey, we conducted a pilot study with three professionals in the field to check the wording and coverage of the proposed answers. Furthermore, we provided “Other” options in the survey, allowing participants to freely input additional reasons and mitigation solutions for ML service misuses. None of the 50 practitioners suggested any additional reasons or mitigation solutions, which indicates that the predefined options comprehensively captured the key factors related to these elements. Despite these efforts, we acknowledge that further research is needed to deepen our understanding of the underlying causes of ML service misuses in software systems and to develop more effective mitigation strategies.
7. Conclusion
This paper presented a comprehensive, multi-vocal empirical study of misuses of ML services. We propose a catalog of 20 ML service misuses based on a systematic literature review, a gray literature review, and an analysis of GitHub projects. We defined these misuses in detail, along with their impact on ML service-based systems and the stage at which they are more likely to occur. We validated our catalog through an online survey with ML practitioners, in which we measured the level of their agreement with our catalog and the frequency of encountering each misuse in practice. In addition, we reported, based on practitioners’ experience, the common reasons behind the presence of such misuses and the recommended solutions to mitigate them. Our study offers developers interested in ML cloud services a thorough understanding of the bad practices to avoid when developing ML service-based systems. Moreover, our study raises awareness of ML service misuses and encourages the involvement of various stakeholders, including cloud providers, open-source platforms, practitioners, and researchers, to mitigate them. In future work, we aim to develop a tool to automatically detect the presence of these misuses, enhancing the accuracy and speed of identification processes. We also intend to develop an automated refactoring solution to effectively remove such misuses, thereby improving code quality and maintainability.